Reports 1-1 of 1 Clear search Modify search
CAL (General)
Shingo hido - 18:06 Wednesday 30 April 2025 (33609) Print this report
CAL rehearsal (FULL calibration)
CAL group

Abstract

We conducted the FULL calibration as CAL rehearsal.
The estimated parameters related to DARM are as follows.
Details about other degrees of freedom will be provided later.

1. Actuator efficiencies of ETMX
H_etmxtm = 3.797563e-14 [m/ct] at 10 Hz (+0.07 % from 4/7 klog#33286)
H_etmxim = 1.677816e-14 [m/ct] at 10 Hz (+ 3 % from 4/7 klog#33286)

2. Optical Gain
Optical gain = 2.492139e12 [ct/m] (- 0.7 % from 4/7 klog#33286)

Details
The actuator efficiencies were estimated using the same method described in klog#32745.
Fig. 1 shows DARM OLTF model + fitting results. Fig. {2, 3} show ratio ETMXTM/PCAL and ETMXIM / ETMXTM.
Measured transfer functions are stored in "/users/CAL/current/measurements/2025/0430/".
Images attached to this report
Comments to this report:
takahiro.yamamoto - 4:22 Thursday 01 May 2025 (33616) Print this report
I summarized all estimated actuator efficiencies as follows.
 |--------------+------------+------------------+--------------------------------+--------------------------------|
| | Prev PCAL | Prev FSM | PCAL | FSM |
|--------------+------------+------------------+--------------------------------+--------------------------------|
| A_etmx_tm | 3.7791e-12 | 3.129e-12 (250K) | 3.8106000e-12 pm 2.7301000e-15 | 3.4851971e-12 pm 7.5046314e-14 |
| A_itmx_tm | | | 3.3965594e-12 pm 3.7797149e-15 | 3.1065132e-12 pm 6.6839839e-14 |
|--------------+------------+------------------+--------------------------------+--------------------------------|
| A_bs_tm_mich | | 6.060e-11 | 6.8162942e-11 pm 4.5444750e-13 | 6.2342227e-11 pm 1.2772218e-12 |
| A_bs_tm_L | | | 4.8198479e-11 pm 3.2134291e-13 | 4.4082612e-11 pm 9.0313223e-13 |
| A_bs_tm_prcl | | | 3.4081471e-11 pm 2.2722375e-13 | 3.1171114e-11 pm 6.3861092e-13 |
|--------------+------------+------------------+--------------------------------+--------------------------------|
| A_prm_tm | | 7.401e-10 | 9.2868600e-10 pm 6.5799986e-12 | 8.4938167e-10 pm 1.7520325e-11 |
|--------------+------------+------------------+--------------------------------+--------------------------------|

- Results by Pcal and FSM are consistent within 9% each other.
- Estimated actuator efficiency of ETMX is no significant changes.
(Previous FSM was done in 250K mirror temperature. So it seems to be a cause of difference.)
- Changes in the PRM actuator efficiency from the past is large. It might be related to the actuator efficiency ratio below.


Actuator efficiency ratio is as follows.
|-----------------------------+-----------+------------------------|
| | prev | |
|-----------------------------+-----------+------------------------|
| itmx_tm_mich / bs_tm_mich | 4.808e-02 | 4.9830e-2 pm 3.2756e-4 |
| etmx_tm_DARM / itmx_tm_DARM | 1.077e-0 | 1.1219e-0 pm 9.5529e-4 |
| prm_tm_PRCL / bs_tm_PRCL | 2.4540e+1 | 2.7249e+1 pm 6.5348e-2 |
|-----------------------------+-----------+------------------------|

Ratio between PRM and BS with PRCL is more than 10% different from the past. I'm suspecting that the modeling of feed-forward effect may be inadequate, though I don't have proof yet. Theoretical calculation will be done tomorrow. And also, re-measuring without feedforward and/or with PRX lock may help us to understand the situation.
takahiro.yamamoto - 11:20 Friday 02 May 2025 (33650) Print this report
I heard that front-end calibration of DARM was updated as new calibration parameters estimated from CAL rehearsal data, but...

Updates of foton file hasn't been loaded yet by the front-end model. So it's sitll working with old values.
Pls. ensure to complete own work...
Images attached to this comment
Shingo Hido - 11:37 Friday 02 May 2025 (33651) Print this report
I'm sorry. I forgot to load the filters.
takahiro.yamamoto - 12:26 Wednesday 07 May 2025 (33695) Print this report

> Ratio between PRM and BS with PRCL is more than 10% different from the past. I'm suspecting that the modeling of feed-forward effect may be inadequate,
Difference in the actuator efficiency ratio between PRM and BS as ~10% cannot be explain the effect of FF.
So there seems to be an another reason.
Re-measurement with PRX might be necessary though I have no good idea about a difference between PRX and (PR)PRMI.

-----
In PRMI (and PRFPMI), MICH and PRCL loops have correlation via BS longitudinal motion as shown in Fig.5 of klog#24980. Actuator efficiency of PRM is estimated based on this fact by using an efficiency ration between BS and PRM. In the case with MICH2PRCL FF, control loops become slightly complicated as shown in Fig.1. We measured actuator efficiency of PRM with FF in klog#33609 and actuator efficiency ratio between PRM and BS was ~27.2. This value is ~11% different from past value (~24.5) in klog#29974 and klog#30091 which is measured with PRX lock. (Because BS and PRM are always in room temperature, I have no good idea about a ~10% change.)

I wondered that this difference comes from the effect of FF and compared the actuator efficiency ratio in PRFPMI with and without FF (measurement files are saved in /users/CAL/current/measurements/2025/0503/. With FF, estimated actuator efficiency ratio is reproduced as ~27.2 (see Fig.2-3). Without FF, it shows ~27.5 which is almost same as a result with FF (see Fig.4-5).​​​​​ So our measurement doesn't seem to be affected by FF.

By the way, as I checked latest foton file, a servo gain is set as 0.03875 in LSC-MICHFF2 which corresponds to ~25.8 as an actuator efficiency ratio between PRM and BS (see also Fig.6). According to the filter name (FFgain231226), it's decided by a measurement in klog#28094. On this klog, a measurement was done as follows.
> 6. Calculate the ratio between MICH2PRCL to PRCLout2PRCL (optimized MICH2PRCL FF).
If "PRCLout" means LSC-PRCL2_OUT, only MICH2PRCL is suppressed by PRCL OLTF though PRCLout2PRCL isn't so because LSC-PRCL2_OUT is in-loop TP. Was this effect corrected? (or negligible due to use only much higher frequency than PRCL UGF?) I couldn't find measurement file so I couldn't check.

Images attached to this comment
Shingo Hido - 12:53 Monday 12 May 2025 (33748) Print this report

In the previous measurement of the MN stage, the measurement at 46.21 Hz failed, so I recreated the frequency list and updated the templates.
The script to generate the frequency list is located at `/users/shingo/CAL/line_freq/cal_line_s.py`.(detailed document in JGW-G16675

Search Help
×

Warning

×